OpenStudy (anonymous):

i dont get this one... a travel agency offers a trip for $550. there is a 10% service fee... if you dont get 30% off the price, how much will you pay for a service fee??

OpenStudy (anonymous):

10/100 = x/550 x = your service fee.

OpenStudy (anonymous):

the next question says you get a 30% off whats the price includeing service of 10%?

OpenStudy (anonymous):

30/100 = x/550. Solve for x. Now we find the new service fee. 10/100 = x/whatever u got for the first part. solve for x. now add your new service fee to what you got for the first part.

OpenStudy (anonymous):

Oh I forgot for the first part subtract whatever you got for x from 550.

OpenStudy (anonymous):

x= 423.5?

OpenStudy (anonymous):

Correct you got it! :D

OpenStudy (anonymous):

and if theres no service fee at all does it cost 55

OpenStudy (anonymous):

*30% off sorry

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

if you bought the $550 trip before the 30% off but including the 10% service..10/100 = x/550... 100x = 5500/100 x=55? is that how i do it or no

OpenStudy (anonymous):

Thats right and you add the service fee to 550

OpenStudy (anonymous):

oh right :P thank youuuuuuu

OpenStudy (anonymous):

no problem man :) I'm gone for the night though. Hope you get help with whatever else you need. There are a lot of people here who can help you

OpenStudy (anonymous):

ya actually that was my last problem thank you sleep nice